A spring has a mass, m =2.5 kg and constant, k = 30 N/m. What is the period of this spring?


1
Expert's answer
2021-01-14T10:38:16-0500

We can find the period of the spring from the formula:


"T=2\pi\sqrt{\dfrac{m}{k}},""T=2\pi\sqrt{\dfrac{2.5\ kg}{30\ \dfrac{N}{m}}}=1.81\ s."

Answer:

"T=1.81\ s."